a) John is now 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages.
Answer by mananth(16946)

Sam's age x
John's age 3x
5 years later
Sam's age x+5
John's age 3x+5
John's age 2 times older
3x+5=2(x+5)
3x+5=2x+10
x=10-5
x=5
x= 5 Sam's age
3x=15 John's age